# Bandicoot: A Templated C++ Library for GPU Linear Algebra

### v1

@@ 1 Introduction
A number of recent significant advancements in scientific computing have been enabled by the use of graphics accelerators / graphics processing units (GPUs), notably in the fields of machine learning [9, 24], astrophysics [16], weather simulations [18], data visualisation [36], as well as healthcare and life sciences [31, 42]. GPUs are built on the concept of single-instruction-multiple-threads (SIMT) and are thus capable of performing the same operation on separate chunks of data in parallel. Often, modern GPUs are extremely parallel: Nvidia’s recent mainstream consumer-focused GeForce RTX 5090 GPU contains 21,760 cores, leading to a throughput of over 100 TFLOPS for 32-bit floating point data.
@@
High-level interfaces have seen much success over the years, specifically as they allow developers and scientists to maximise productivity by focusing on their higher-level task. In this vein, frameworks like MATLAB and Mathematica [47] have been extremely successful, in large part because of the simple and intuitive interface they provide. In the field of machine learning, Python [43] and its linear algebra abstraction NumPy [17] have become dominant in large part for the same reason. However, these higher-level languages often come with various downsides; for instance, MATLAB and Python have high amounts of overhead, and require large runtime environments that depend on numerous additional packages.
Motivated by high-level MATLAB-style expressive syntax, the well-known Armadillo C++ linear algebra library was created to satisfy the need for a high-level and intuitive linear algebra interface while maximising runtime performance and minimising overheads [32, 33]. Armadillo extensively uses template meta-programming [44] in order to optimise linear algebra expressions at program compilation time; this can provide significant and demonstrable speedups over other solutions [29]. Its clear interface and flexibility has led to widespread adoption, including via bindings to other languages like R [13] and Python [41].
@@ 2 Overview of Functionality
At the time of writing, Bandicoot provides over 100 functions and is comprised of about 85,000 lines of templated C++ code. In addition to elementary arithmetic operations such as matrix addition and multiplication, there are functions for statistics (such as mean and variance), signal processing (such as 1D and 2D convolution), diagonal and submatrix views (allowing operations to use and be applied to parts of matrices), broadcasting operations, various matrix factorisations and decompositions (including Cholesky decomposition, eigen decomposition of symmetric matrices, lower-upper (LU) decomposition, singular value decomposition), matrix inverses, and a solver for system of linear equations. Wherever possible, the provided functions aim to be as similar as possible to Matlab, closely following the approach used by Armadillo [33]. A comprehensive documentation of all available functionality in Bandicoot is provided online at https://coot.sourceforge.io/docs.html.
@@ 3 Internal Design
All matrix and vector objects as well as all linear algebra expressions inherit from the fundamental class Base<eT,T1>, where eT refers to the element type and T1 refers to a specific class representing a matrix, vector, or compound linear algebra expression. This is a form of static polymorphism (polymorphism without runtime virtual-table lookups) modelled after the Curiously Recurring Template Pattern [3, 10]. User-facing classes include the Mat<eT>, Col<eT>, and Row<eT> classes, which represent matrices, column vectors, and row vectors, respectively. The vector classes are derived from the matrix class, which allows the treatment of vectors as specialised versions of matrices. A number of predefined short-hand typedefs are given, allowing users to use more concise forms such as fmat type instead of Mat<float>.
@@ 1. $\bullet$
    The primary goal for Bandicoot is to be API-compatible with Armadillo as much as possible, to ease adaptation of Armadillo based programs. To that end, we are currently in the process of expanding Bandicoot’s API with additional matrix decompositions, such as QR, QZ, Schur and generalised eigenvalues [14, 39], as well as expansion of the solve() function to match the multitude of options present in Armadillo. Furthermore, we are adding support for complex numbers, to allow matrices such as Mat< std::complex<float> >.
@@ 2. $\bullet$
    Use of CLBlast instead of clBLAS. The clBLAS project [1] has not been updated in years and has bugs that are likely to remain unfixed. The newer CLBlast project provides an appealing higher-performance alternative implementation of BLAS functions on OpenCL [27]. Like Armadillo’s support for any BLAS/LAPACK replacement, the goal for Bandicoot is to allow the use of either library with the OpenCL backend.

### v2

@@ 1 Introduction
A number of recent significant advancements in scientific computing have been enabled by the use of graphics accelerators / graphics processing units (GPUs), notably in the fields of machine learning [9, 24], astrophysics [16], weather simulations [6, 18], data visualisation [37], as well as healthcare and life sciences [31, 43]. GPUs are built on the concept of single-instruction-multiple-threads (SIMT) and are thus capable of performing the same operation on separate chunks of data in parallel. Often, modern GPUs are extremely parallel: Nvidia’s recent mainstream consumer-focused GeForce RTX 5090 GPU contains 21,760 cores, leading to a throughput of over 100 TFLOPS for 32-bit floating point data.
@@
High-level interfaces have seen much success over the years, specifically as they allow developers and scientists to maximise productivity by focusing on their higher-level task. In this vein, frameworks like MATLAB and Mathematica [47] have been extremely successful, in large part because of the simple and intuitive interface they provide. In the field of machine learning, Python and its NumPy linear algebra add-on [17] have become dominant in large part for the same reason. However, these higher-level languages often come with various downsides; for instance, MATLAB and Python have high amounts of overhead, and require large runtime environments that depend on numerous additional packages.
Motivated by high-level MATLAB-style expressive syntax, the well-known Armadillo C++ linear algebra library was created to satisfy the need for a high-level and intuitive linear algebra interface while maximising runtime performance and minimising overheads [33, 34]. Armadillo extensively uses template meta-programming [44] in order to optimise linear algebra expressions at program compilation time; this can provide significant and demonstrable speedups over other solutions [29]. Its clear interface and flexibility has led to widespread adoption, including via bindings to other languages like R [13] and Python [32, 42].
@@ 2 Overview of Functionality
At the time of writing, Bandicoot provides over 100 functions and is comprised of about 85,000 lines of templated C++ code. In addition to elementary arithmetic operations such as matrix addition and multiplication, there are functions for statistics (such as mean and variance), signal processing (such as 1D and 2D convolution), diagonal and submatrix views (allowing operations to use and be applied to parts of matrices), broadcasting operations, various matrix factorisations and decompositions (including Cholesky decomposition, eigen decomposition, lower-upper (LU) decomposition, singular value decomposition), matrix inverses (including pseudo-inverse), and a solver for system of linear equations. Wherever possible, the provided functions aim to be as similar as possible to MATLAB, closely following the approach used by Armadillo [34]. A comprehensive documentation of all available functionality in Bandicoot is provided online at https://coot.sourceforge.io/docs.html.
@@ 3 Internal Design
All matrix and vector objects as well as all linear algebra expressions inherit from the fundamental class Base<eT,T1>, where eT refers to the element type and T1 refers to a specific class representing a matrix, vector, or compound linear algebra expression. This is a form of static polymorphism, omitting runtime virtual-table lookups [2, 44]. User-facing classes include the Mat<eT>, Col<eT>, and Row<eT> classes, which represent matrices, column vectors, and row vectors, respectively. The vector classes are derived from the matrix class, which allows the treatment of vectors as specialised versions of matrices. A number of predefined short-hand typedefs are given, allowing users to use more concise forms such as fmat type instead of Mat<float>.
@@ 1. $\bullet$
    API expansion. The primary goal for Bandicoot is to be API-compatible with Armadillo as much as possible, to ease adaptation of Armadillo based programs. To that end, we are currently in the process of expanding Bandicoot’s API with additional matrix decompositions, such as QR, QZ, Schur and generalised eigenvalues [14, 40], as well as expansion of the solve() function to match the multitude of options present in Armadillo. Furthermore, we are adding support for complex numbers, to allow matrices such as Mat< std::complex<float> >.
@@ 2. $\bullet$
    Use of CLBlast instead of clBLAS. The clBLAS project [10] has not been updated in many years and has bugs that are likely to remain unfixed. The newer and better maintained CLBlast project provides an appealing higher-performance alternative implementation of BLAS functions on OpenCL [27]. Like Armadillo’s support for any BLAS/LAPACK replacement, the goal for Bandicoot is to allow the use of either library with the OpenCL backend.